The sum of two integers is 36, the difference is 6. What is the smaller of the two numbers?
Answer by Nielsen(1)

x+y=36
+ x-y=6
_______
2x=42
2x/2=42/2
(x=21)
x+y=36
21+y=36
21-21+y=36-21
(y=15)
the answers are 15 and 21
the smaller is 15